# One Token Can Help! Learning Scalable and Pluggable Virtual Tokens for Retrieval-Augmented Large Language Models

### v2

@@ Abstract
Retrieval-augmented generation (RAG) is a promising way to improve large language models (LLMs) for generating more factual, accurate, and up-to-date content. Existing methods either optimize prompts to guide LLMs in leveraging retrieved information or directly fine-tune the LLMs to adapt to RAG scenarios. Although fine-tuning can yield better performance, it often compromises the LLMs’ general generation capabilities by modifying their parameters. This limitation poses challenges in practical applications, especially when LLMs are already deployed, as parameter adjustments may affect their original functionality. To address this, we propose a novel method that involves learning scalable and pluggable virtual tokens for RAG. By maintaining the LLMs’ original parameters and fine-tuning only the embeddings of these pluggable tokens, our approach not only enhances LLMs’ performance but also preserves their general generation capacities. Furthermore, we design several training strategies to improve the scalability, flexibility, and generalizability of our method. Comprehensive experiments across nine question-answering tasks demonstrate the superiority of our approach. The virtual tokens and training code are publicly available at https://github.com/DaoD/SPRING/.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance across various natural language processing tasks Brown et al. (2020); OpenAI (2023); Touvron et al. (2023). Despite their extensive parameters enabling them to learn rich knowledge during pre-training, LLMs may still generate hallucinated, outdated, or inaccurate content, especially in scenarios requiring long-tail knowledge Ji et al. (2023); Zhang et al. (2023b). To address this problem, retrieval-augmented generation (RAG) has emerged as a pivotal strategy. By explicitly decoupling knowledge retrieval from the backbone LLMs, such architectures have achieved more accurate and reliable content generation and shown particularly enhanced performance on knowledge-intensive tasks such as open-domain question answering Lewis et al. (2020b); Izacard et al. (2023); Petroni et al. (2021); Tan et al. (2024); Jin et al. (2024).
@@
Therefore, a critical research problem arises: Is it possible to enhance LLMs’ performance under RAG scenarios while preserving their general generation capabilities? To achieve this, we introduce a novel, lightweight tuning method named SPRING, which learns Scalable and Pluggable viRtual tokens for retrIeval-augmeNted Generation. Our basic idea is to add trainable virtual tokens to help LLMs learn RAG problems. Through fine-tuning, these virtual tokens effectively enhance the LLM’s capacity to understand retrieved information and its correlation with user inputs. Importantly, as the LLM’s original parameters are frozen, its general generation abilities are preserved without any loss. During inference, when retrieval is triggered, these trained virtual tokens can be simply added to the prompt, which includes both the retrieved results and user input, thereby significantly enhancing performance. Moreover, we employ a scalable training approach, allowing the number of virtual tokens to be adjusted according to the needs of the inference scenario. Various training strategies have been implemented to further improve the generalizability of our method, ensuring robustness regardless of the number of the retrieved results.
In experiments, SPRING is trained with Mistral-7b, LLaMA-2-7b, Phi-3-4b, and QWen-1.8b on nine commonly used QA datasets and evaluated in both in-domain and out-of-domain tasks. The experimental results demonstrate that SPRING not only effectively improves the RAG performance of LLMs but also successfully preserves their general generation capabilities. Overall, the SPRING method exhibits four main characteristics:
@@ 3 Methodology
To take advantage of both the flexibility of prompt-based methods and the efficacy of fine-tuning-based methods, we propose SPRING to learn scalable and pluggable virtual tokens for retrieval-augmented generation (RAG). In this section, we will first briefly formulate the problem of RAG (§3.1), and then dive into the details of our method (§3.2).
@@ 3.2 SPRING: Scalable and Pluggable Virtual Tokens for RAG
Our SPRING method, shown in Figure 2, introduces trainable virtual tokens into the input to optimize LLMs for RAG scenarios. Specifically, following the notation in Equation (3), we add $n$ trainable tokens $T=[t_{1},t_{2},\cdots,t_{n}]$ between the retrieved results ${R}_{\mathcal{D}}(Q)$ and the input $Q$. The generation process can then be described as:
@@
Importantly, we place the virtual tokens $T$ between the retrieved results $R$ and the question $Q$ for two main reasons: (1) In the auto-regressive generation paradigm, positioning these tokens after the retrieved results allows them to attend to this information, thereby aiding the model’s comprehension. (2) Recent studies have indicated that LLMs are particularly sensitive to the end of an input sequence Liu et al. (2023). By consistently placing these virtual tokens before the question across all test samples, we aim to mitigate any potential adverse effects on the understanding of the question.
@@
Due to its designed structure, our method provides considerable flexibility in application. Practically, if user input is assessed to require external knowledge, our virtual tokens can be simply appended after the retrieval results and then fed, along with the user input, into the LLM for generation. In contrast, if the user input does not necessitate retrieval, it can be processed directly by the LLM. As our approach does not adjust the original parameters of the LLM, it preserves the the model’s inherent capabilities. This feature is particularly important for industry or business since existing LLMs may have already been deployed for multiple purposes; our method enhances the retrieval understanding capabilities of these models without compromising their existing functionalities.
@@
We illustrate the instructions for using our SPRING in Figure 3. After training, the embeddings of the added tokens have been optimized for RAG, but these tokens do not correspond to any existing tokens in the vocabulary. To make them easy to use, we can add some special tokens (e.g., [r1], $\cdots$, [r50]) to the vocabulary and initialize their embeddings with the trained embeddings. Then, during inference, after obtaining the retrieved results $R$, we can add any number of these special tokens (e.g., [r1] $\cdots$ [rk]) after $R$ and input them with the question $Q$ to the LLMs for generation.
@@
Table 1: Evaluation results of different backbone models and methods on nine QA datasets. The retriever is e5-large model, and the number of retrieved passages is set as three. PopQA is invisible during training. “Prefix” stands for prefix-tuning. The best results are in bold.
@@ 4.1 Datasets and Retrievers
We conduct experiments on nine commonly used question-answering datasets, including TriviaQA (TQA) Joshi et al. (2017), Natural Questions (NQ) Kwiatkowski et al. (2019), HotpotQA (HQA) Yang et al. (2018), SQuAD 1.0 Rajpurkar et al. (2016), Web Questions (WebQ) Berant et al. (2013), 2WikiMultiHopQA (2Wiki) Ho et al. (2020), CoQA Reddy et al. (2019), MS MARCO Nguyen et al. (2016), and PopQA Mallen et al. (2023). These datasets are publicly available at HuggingFace or their official websites. Since PopQA only has a test set, we use it as a held-out dataset to evaluate the generalizability of the methods. Except for PopQA, we mix the training set of all other datasets for training, and test the methods on their corresponding test set. If the test set is unavailable, we use the development set for evaluation. It is worth noting that, though some datasets have provided golden reference passages for the answer, we do not use them in our experiment but use the passages retrieved from the following retrieval sets in both training and inference stages. Exact match (EM) and F1 score are employed as evaluation metrics.
For the retrieval sets, we follow previous studies Yoran et al. (2023) and use the combination of Wikipedia and MS MARCO datasets as the retrieval corpus. Wikipedia contains high-quality human knowledge, which is helpful for many knowledge-intensive tasks. MS MARCO contains a large amount of web pages, which can provide information necessary for curating some natural language questions. We use the datasets that have already been preprossed into passages and released on HuggingFace. [^1] Wikipedia passages: https://huggingface.co/datasets/Tevatron/wikipedia-nq-corpus. MS MARCO passages: https://huggingface.co/datasets/Tevatron/msmarco-passage-corpus. The Wikipedia set has 21M passages, while the MS MARCO set has 8M passages. More details will be provided in Appendix.
We use E5-large Wang et al. (2022) as the main retriever in our experiments. The impact of other retrievers, i.e., BM25 Robertson and Zaragoza (2009), BGE-base Xiao et al. (2023), and E5-base Wang et al. (2022), are studied in our further analysis. Among these retrievers, BM25 is a non-neural sparse retrieval algorithm, while others are neural-based dense retrievers. In general, dense retrievers perform better than BM25 on several benchmarks Wang et al. (2022); Izacard et al. (2022).
@@ 4.2 Baseline Methods
Mistral-7b-Instruct, LLaMA-2-7b-Chat Touvron et al. (2023), Phi-3 Mini-4K-Instruct Abdin et al. (2024), and Qwen1.5-4B-Chat Bai et al. (2023) are used as backbone models. We compare our method with the following baselines.
@@
$\bullet$ Prompt: This method uses a manually-crafted prompt to indicate the use of retrieval information (details will be added in Appendix).
@@ 4.3 Experimental Results
We fine-tune the prefix-tuning, LoRA, and SPRING methods on RAG tasks, and then evaluate their performance in scenarios both with (RAG) and without (non-RAG) retrieval. The experimental results are shown in Table 1. To save space, we only show the results based on Mistral-7b-Instruct, and other results will be available at Appendix.
We can observe: (1) It is evident that SPRING significantly improves the RAG performance of the original LLM with manually-crafted prompts (the average EM and F1 scores are improved by $33.9\%$ and $12.7\%$, respectively). It even outperforms LoRA on certain datasets, such as TriviaQA and CoQA. Given that SPRING involves only 0.2M trainable parameters, these results demonstrate its remarkable efficiency and effectiveness. (2) While LoRA achieves slightly better performance, it adjusts the LLMs’ original parameters, which adversely impact their performance in non-RAG scenarios — a significant drop has been observed, far worse than the original models. This challenge also extends to other general generation tasks, which will be discussed in the next section. (3) All backbone models show improvements with our SPRING, demonstrating its adaptability across various backbone models. This adaptability is further validated by consistent performance improvements for both the instruct-tuned and base models (will be added in Appendix). (4) SPRING achieves the best performance on the held-out test set, PopQA, validating the good generalizability of our method. (5) Interestingly, prefix-tuning cannot perform well for RAG, highlighting that the insertion position of the virtual tokens in SPRING is both reasonable and effective.
@@ 4.4 Further Analysis
Table 2: Performance of LoRA and SPRING on other general tasks.
@@ 4.4.1 Performance on Other Tasks
To examine the impact of different fine-tuning methods on the inherent capabilities of LLMs, we evaluate the performance of models fine-tuned by LoRA and SPRING on several other (non-RAG) tasks. These tasks are commonly used to evaluate LLMs’ reasoning, mathematical abilities, and world knowledge, including BoolQ Clark et al. (2019), CommonsenseQA Talmor et al. (2019), GSM8K Cobbe et al. (2021), and MMLU Hendrycks et al. (2021). The experimental results are shown in Table 2.333We notice that our results are quite different from those officially reported, which we attribute to the impact of different prompts. Since official prompts for testing are unavailable, we provide the prompts we used in Appendix LABEL: to facilitate reproducibility. From the results, we can observe: (1) Thanks to the plug-and-play design of our method, SPRING can revert to to the original LLMs by not using virtual tokens. Therefore, it successfully preserves the original capabilities of the LLMs. In contrast, LoRA, which adjusts the model’s parameters for RAG tasks, inevitably compromises the model’s performance on other tasks. (2) A noticeable decline is observed in the few-shot evaluation, reflecting a decrease in the in-context learning abilities of LLMs. This decline may stem from the fact that RAG fine-tuning does not incorporate in-context learning capabilities. Besides, fine-tuning for RAG tasks may lead the model to overfit to specific task formats, thereby impairing its general generation abilities.
@@ 4.4.2 Impact of Token Quantity
In SPRING, we design a scalable training approach that enables to use arbitrary numbers of virtual tokens in inference. To validate its effectiveness, we test the performance of our method with various numbers of virtual tokens and compare it with a variant model trained with a fixed number of tokens ($n=50$). The experimental results are illustrated in Figure 4. In general, we can observe that the performance of SPRING increases with more virtual tokens used. Surprisingly, SPRING can significantly enhance LLMs’ performance in RAG scenarios with just a single token, which is very encouraging. In comparison, training with a fixed number of tokens limits the flexibility of SPRING, making it can only be used with the same number of tokens in inference (i.e., $n=50$).
@@ 4.4.4 Influence of Retrieved Passages
In the fine-tuning process of SPRING, we construct training samples by randomly selecting the top-$m$ ($m\in[1,5]$) retrieved passages. This strategy aims to enhance SPRING’s adaptability by ensuring it can operate effectively with varying numbers of retrieved passages in real-world scenarios. To evaluate the effect of this training strategy, we test the SPRING’s performance across a range from zero to five passages. Figure 5 illustrates the results. We can find that SPRING’s performance gradually improves as more retrieved passages are used ($m=0\rightarrow 4$), suggesting that more retrieved passages contribute valuable knowledge for question answering. However, the performance peaks at four passages and declines when more passages are added. This decrease could be attributed to noise accumulation within the retrieved knowledge, a phenomenon also reported in recent studies Yoran et al. (2023); Wu et al. (2024). Despite this, the use of retrieved passages still results in performance gains compared to scenarios without retrieval ($m=0$), highlighting again the benefits of RAG.

### v3

@@ Abstract
Retrieval-augmented generation (RAG) is a promising way to improve large language models (LLMs) for generating more factual, accurate, and up-to-date content. Existing methods either optimize prompts to guide LLMs in leveraging retrieved information or directly fine-tune LLMs to adapt to RAG scenarios. Although fine-tuning can yield better performance, it often compromises the LLMs’ general generation capabilities by modifying their parameters. This limitation poses challenges in practical applications, especially when LLMs are already deployed, as parameter adjustments may affect their original functionality. To address this, we propose a novel method that involves learning scalable and pluggable virtual tokens for RAG. By maintaining the LLMs’ original parameters and fine-tuning only the embeddings of these pluggable tokens, our approach not only enhances LLMs’ performance but also preserves their general generation capabilities. Furthermore, we design several training strategies to improve the scalability, flexibility, and generalizability of our method. Comprehensive experiments across nine question-answering tasks demonstrate the superiority of our approach. The virtual tokens and training code are publicly available at https://github.com/DaoD/SPRING/.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance across various natural language processing tasks Brown et al. (2020); OpenAI (2023); Touvron et al. (2023). Despite their extensive parameters enabling them to learn rich knowledge during pre-training, LLMs may still generate hallucinated, outdated, or inaccurate content, especially in scenarios requiring long-tail knowledge Ji et al. (2023); Zhang et al. (2023b). To address this problem, retrieval-augmented generation (RAG) has emerged as a pivotal strategy. By explicitly decoupling knowledge retrieval from the backbone LLMs, such architectures have achieved more accurate and reliable content generation and shown particularly enhanced performance on knowledge-intensive tasks such as open-domain question answering Lewis et al. (2020b); Petroni et al. (2021); Izacard et al. (2023); Tan et al. (2024); Jin et al. (2024).
@@
Therefore, a critical research problem arises: Is it possible to enhance LLMs’ performance under RAG scenarios while preserving their general generation capabilities? To achieve this, we introduce a novel, lightweight tuning method named SPRING, which learns Scalable and Pluggable viRtual tokens for retrIeval-augmeNted Generation. Our basic idea is to add trainable virtual tokens to help LLMs learn RAG problems. Through fine-tuning, these virtual tokens effectively enhance the LLM’s capability to understand retrieved information and its correlation with user inputs. Importantly, as the LLM’s original parameters are frozen, its general generation abilities are preserved without any loss. During inference, when retrieval is triggered, these trained virtual tokens can be simply added to the prompt, which includes both the retrieved results and user input, thereby significantly enhancing performance. Moreover, we employ a scalable training approach, allowing the number of virtual tokens to be adjusted according to the needs of the inference scenario. Various training strategies have been implemented to further improve the generalizability of our method, ensuring robustness regardless of the number of the retrieved results.
In experiments, SPRING is trained with Mistral-7b, LLaMA-2-7b, and LLaMA-2-13b on nine commonly used QA datasets and evaluated in both in-domain and out-of-domain tasks. The experimental results demonstrate that SPRING not only effectively improves the RAG performance of LLMs but also successfully preserves their general generation capabilities. Overall, the SPRING method exhibits four main characteristics:
@@ 3 Methodology
To take advantage of both the flexibility of prompt-based methods and the efficacy of fine-tuning-based methods, we propose SPRING to learn scalable and pluggable virtual tokens for retrieval-augmented generation (RAG).
@@ 3.2 SPRING: Scalable and Pluggable Virtual Tokens for RAG
Our SPRING method, shown in Figure 2, introduces trainable virtual tokens into the input to optimize LLMs for RAG scenarios. Specifically, following the notation in Equation (3), we add $n$ trainable tokens $T=[t_{1},t_{2},\cdots,t_{n}]$ between the retrieved results ${R}_{\mathcal{D}}(Q)$ and the input $Q$ [^1] In our implementation, we set $n=50$ for training. The generation process can then be described as:
@@
Importantly, we place the virtual tokens $T$ between the retrieved results $R$ and the question $Q$ for two main reasons: (1) In the auto-regressive generation paradigm, positioning the tokens after the retrieved results allows them to attend to this information, thereby aiding the model’s comprehension. (2) Recent studies have indicated that LLMs are particularly sensitive to the end of an input Liu et al. (2023). By consistently placing these virtual tokens before the question across all test samples, we aim to mitigate any potential adverse effects on the understanding of the question.
@@
Due to its designed structure, our method provides considerable flexibility in application. Practically, if user input is assessed to require external knowledge, our virtual tokens can be simply appended after the retrieval results and then fed, along with the user input, into the LLM for generation. In contrast, if the user input does not necessitate retrieval, it can be processed directly by the LLM. As our approach does not adjust the original parameters of the LLM, it preserves the model’s inherent capabilities. This feature is particularly important for industry or business since existing LLMs may have already been deployed for multiple purposes; our method enhances the retrieval understanding capabilities of these models without compromising their existing functionalities.
@@
We illustrate the instructions for using our SPRING in Figure 3. After training, the embeddings of the added tokens have been optimized for RAG, but these tokens do not correspond to any existing tokens in the vocabulary. To make them easy to use, we can add some special tokens (e.g., [r1], $\cdots$, [r50]) to the vocabulary and initialize their embeddings with the trained embeddings. Then, during inference, after obtaining the retrieved results $R$, we can add any number of these special tokens (e.g., [r1] $\cdots$ [rk]) after $R$ and input them with the question $Q$ to the LLMs for generation. We also provide an example code snippet in Appendix A for using our method in practice.
@@
Table 1: Evaluation results of different backbone models and methods on nine QA datasets. The retriever is e5-large model, and the number of retrieved passages is set as three. The number of virtual tokens used in SPRING is set as 50. PopQA is invisible during training. “Prefix” stands for prefix-tuning. The best results are in bold.
@@ 4.1 Datasets and Retrievers
We conduct experiments on nine commonly used question-answering datasets, including TriviaQA (TQA) Joshi et al. (2017), Natural Questions (NQ) Kwiatkowski et al. (2019), HotpotQA (HQA) Yang et al. (2018), SQuAD 1.1 Rajpurkar et al. (2016), Web Questions (WebQ) Berant et al. (2013), 2WikiMultiHopQA (2Wiki) Ho et al. (2020), CoQA Reddy et al. (2019), MS MARCO Nguyen et al. (2016), and PopQA Mallen et al. (2023). These datasets are publicly available at HuggingFace or their official websites. Since PopQA only has a test set, we use it as a held-out dataset to evaluate the generalizability of the methods. Except for PopQA, we mix the training set of all other datasets for training, and test the methods on their corresponding test set. If the test set is unavailable, we use the development set for evaluation. It is worth noting that, though some datasets have provided golden reference passages for the answer, we do not use them in our experiment but use the passages retrieved from the following retrieval sets in both training and inference stages. Exact match (EM) and F1 score are employed as evaluation metrics.
For the retrieval sets, we follow previous studies Yoran et al. (2023) and use the combination of Wikipedia and MS MARCO datasets as the retrieval corpus. Wikipedia contains high-quality human knowledge, which is helpful for many knowledge-intensive tasks. MS MARCO contains a large amount of web pages, which can provide information necessary for curating some natural language questions. We use the datasets that have already been preprossed into passages and released on HuggingFace. [^2] Wikipedia passages: https://huggingface.co/datasets/Tevatron/wikipedia-nq-corpus. MS MARCO passages: https://huggingface.co/datasets/Tevatron/msmarco-passage-corpus. The Wikipedia set has 21M passages, while the MS MARCO set has 8M passages. More details are provided in Appendix B.
We use E5-large Wang et al. (2022) as the main retriever in our experiments. The impact of other retrievers, i.e., BM25 Robertson and Zaragoza (2009), BGE-base Xiao et al. (2023), and E5-base, is studied in our further analysis (Section 4.4.3). Among these retrievers, BM25 is a non-neural sparse retrieval algorithm, while others are neural-based dense retrievers. In general, dense retrievers perform better than BM25 on several benchmarks Wang et al. (2022); Izacard et al. (2022).
@@ 4.2 Baseline Methods
We consider both the base and instruction-tuned versions of Mistral-7b, LLaMA-2-7b, and LLaMA-2-13b as the backbone models, and compare our SPRING with the following baselines.
@@
$\bullet$ Prompt: This method uses a manually-crafted prompt to indicate the use of retrieval information (details are provided in Appendix C).
@@ 4.3 Experimental Results
We fine-tune the prefix-tuning, LoRA, and SPRING methods on RAG tasks, and then evaluate their performance in scenarios both with (RAG) and without (non-RAG) retrieval. For SPRING, we use $k=50$ virtual tokens for inference by default, and the impact of token quantity $k$ is discussed in Section 4.4.2. The experimental results are shown in Table 1. To save space, we only show the results based on Mistral-7b-instruct, and other results are provided in Appendix E.
We can observe: (1) It is evident that SPRING significantly improves the RAG performance of the original LLM with manually-crafted prompts (the average EM and F1 scores are improved by $33.9\%$ and $12.7\%$, respectively). It even outperforms LoRA on certain datasets, such as TriviaQA and CoQA. Given that SPRING involves only 0.2M trainable parameters, these results demonstrate its remarkable efficiency and effectiveness. (2) While LoRA achieves slightly better performance, it adjusts the LLMs’ original parameters, which adversely impact their performance in non-RAG scenarios — a significant drop has been observed, far worse than the original models. This challenge also extends to other general generation tasks, which will be discussed in the next section. (3) Using manually-crafted prompts is effective for improving LLMs’ performance on RAG tasks. However, this improvement is limited as no training is involved. (4) All backbone models show improvements with SPRING (see Appendix E), demonstrating its adaptability across various backbone models. (5) SPRING achieves the best performance on the held-out test set, PopQA, validating the good generalizability of our method. (6) Interestingly, prefix-tuning cannot perform well for RAG, highlighting that the insertion position of the virtual tokens in SPRING is both reasonable and effective. We conduct more analysis in Appendix F.
@@ 4.4 Further Analysis
Table 2: Performance comparison on other datasets.
@@ 4.4.1 Performance on Other Tasks
To examine the impact of different fine-tuning methods on the inherent capabilities of LLMs, we evaluate the performance of models fine-tuned by LoRA and SPRING on several other (non-RAG) tasks. These tasks are commonly used to evaluate LLMs’ reasoning, mathematical abilities, and world knowledge, including BoolQ Clark et al. (2019), CommonsenseQA Talmor et al. (2019), GSM8K Cobbe et al. (2021), and MMLU Hendrycks et al. (2021). The experimental results are shown in Table 2.444We notice that our results are quite different from those officially reported, which we attribute to the impact of different prompts. Since official prompts for testing are unavailable, we provide the prompts we used in Appendix C to facilitate reproducibility. From the results, we can observe: (1) Thanks to the plug-and-play design of our method, SPRING can revert to to the original LLMs by not using virtual tokens. Therefore, it successfully preserves the original capabilities of the LLMs. In contrast, LoRA, which adjusts the model’s parameters for RAG tasks, inevitably compromises the model’s performance on other tasks. (2) A noticeable decline is observed in the few-shot evaluation, reflecting a decrease in the in-context learning abilities of LLMs. This decline may stem from the fact that RAG fine-tuning does not incorporate in-context learning capabilities. Besides, fine-tuning for RAG tasks may lead the model to overfit to specific task formats, thereby impairing its general generation abilities (more empirical studies are detailed in Appendix G).
@@ 4.4.2 Impact of Token Quantity
In SPRING, we design a scalable training approach that enables to use arbitrary numbers of virtual tokens in inference. To validate its effectiveness, we test the performance of our method with various numbers of virtual tokens and compare it with a variant model trained with a fixed number of tokens ($k=50$). The experimental results are illustrated in Figure 4. In general, we can observe that the performance of SPRING increases with more virtual tokens used. Surprisingly, SPRING can significantly enhance LLMs’ performance in RAG scenarios with just a single token, which is very encouraging. [^5] This varies across different LLMs (see Appendix H). In comparison, training with a fixed number of tokens limits the flexibility of SPRING, making it can only be used with the same number of tokens in inference (i.e., $k=50$).
@@ 4.4.4 Influence of Retrieved Passages
During the fine-tuning of SPRING, we construct training samples by randomly selecting the top-$m$ ($m\in[1,5]$) retrieved passages. This aims to enhance SPRING’s adaptability by ensuring it can operate effectively with varying numbers of retrieved passages in real-world scenarios. To evaluate the effect of this training strategy, we test the SPRING’s performance across a range from zero to five passages. Figure 5 illustrates the results. We can find that SPRING’s performance gradually improves as more retrieved passages are used ($m=0\rightarrow 4$), suggesting that more retrieved passages contribute valuable knowledge for question answering. However, the performance peaks at four passages and declines when more passages are added. This decrease could be attributed to noise accumulation within the retrieved knowledge, a phenomenon also reported in recent studies Yoran et al. (2023); Wu et al. (2024). Despite this, the use of retrieved passages still results in performance gains compared to scenarios without retrieval ($m=0$), highlighting again the benefits of RAG.